# Exploring Multiple Strategies to Improve Multilingual Coreference Resolution in CorefUD

### v2

@@ 3.3 Evaluation Metrics
For evaluation, we employ the official metric for CRAC 2023 Multilingual Coreference Resolution shared task implemented via the CoreUD scorer222https://github.com/ufal/corefud-scorer. The metric is a modification of the CoNLL score for coreference resolution, averaging three metrics: CEAF-e, , and MUC. All these three metrics are described later in this section. The main modification of the CorefUD scorer lies in using the head matching. In head matching the gold and system mentions are considered identical if and only if they have the same syntactic head. Another important aspect is that the primary metric ignores singletons. The CorefUD scorer also provides additional metrics, such as BLANC [20] and LEA [21]. To assess the quality of mention matching while ignoring the assignment of mentions to coreferential entities, it uses the MOR score (mention overlap ratio) [13].
3.3.1 A Link-Based Metric: The MUC Score
The MUC official scorer [22] introduced a link-based metric. A link-based metric measures the extent to which the links in the response match the links in the key. For example, recall is computed by summing up the correctly recalled links for each coreference chain in the key and then dividing by the total number of correct links in the key. The number of missing links—the links found in the key entities but not in the response entities—is computed by counting the number of partitions of key K induced by response R, as follows:
@@ 3.3.1 A Link-Based Metric: The MUC Score
where is the partition function, which returns all the partitions of key entity with respect to a system’s response . Coreference chains are formed by linking each mention in a cluster to its antecedent, creating a sequence of references. Precision is computed by summing up the correct links in each coreference chain in the response and dividing that by the total number of links in the response. Essentially, precision evaluates how many of the predicted coreference links accurately match the gold standard. This is done by swapping key and response in the formula, where the system’s response is compared against the gold coreference chain, and the number of correct links is divided by the total predicted links.
3.3.2 A Mention-Based Metric:
One problem with the MUC score is that, by definition, it only scores a system’s ability to identify links between mentions; its ability to recognize that a mention does not belong to any coreference chain—that is, its ability to classify a mention as a singleton—does not get any reward. The metric [23] was proposed to correct this problem. It does this by computing recall and precision for each mention m, even if m is a singleton. computes the intersection between every coreference chain in the key and every coreference chain in the response and then sums up recall and precision for each pair , and normalizes. In turn, recall and precision for , are computed by summing up recall and precision for each mention m in . For instance, recall for is the proportion of mentions in and the number of mentions in :
Precision for is the proportion between and .
3.3.3 An Entity-Based Metric: CEAF
also suffers from a problem—namely, that a single chain in the key or response can be credited several times. This leads to anomalies; for instance, if all coreference chains in the key are merged into one in the response, the recall is one. The ceaf metric was proposed by [24] to correct this problem. The key idea of CEAF is to align chains (entities) in the key and response using a map in such a way that each chain in the key is aligned with only one chain in the response and to then use the similarity to compute recall and precision. Because different maps are possible, the one that achieves optimal similarity is used. [24] present two entity-pair similarity measures that can be used in CEAF: one is the absolute number of common mentions between two entities, and the other is a “local” mention F-measure between two entities. The two measures lead to the mention-based CEAF-m and entity-based CEAF-e, respectively. [25]
@@ 4 Baseline Model & Extensions
We adopt a model from [1] and we use it as the baseline model (without higher-order inference). We use UDAPI333https://github.com/udapi/udapi-python to load the data in CorefUD format. We use XLM-Roberta-large as an encoder.
Building on this baseline, we propose several extensions to the standard end-to-end model presented earlier in this paper. Our objective is to create a universal model suitable for all datasets in the CorefUD collection.
@@ 4.1 Cross-lingual Training
mBERT | 180M | 40M |
XLM-R | 550M | 50M |
As you can see from Table 4, there are approximately 50 million parameters trained from scratch for XLM-R. For smaller datasets, it is practically impossible to train so many random parameters from scratch. To address this issue, we first pre-train the model on the joined dataset and then fine-tune the model for a specific language.
@@ 4.2 Syntactic Information
We believe that incorporating syntactic information can significantly enhance the model, particularly when manually annotated dependencies are available, such as in the Czech PDT dataset. Moreover, dependency information is essential for identifying mention heads, as some datasets in the CorefUD collection annotate mentions as nodes in the dependency tree rather than spans in the linear text flow.
To encode syntactic information, we add to each token representation its path to the ROOT in the dependency tree. In more detail, we first set the maximum tree depth parameter and then concatenate Bert representations of all parents up to max depth with the embedding of the corresponding dependency relation. Thus the resulting tree structure representation has the size of . This representation is then concatenated with BERT embedding of each token.
@@ 4.3 Span2head Model
Given that the official scorer uses head-match evaluation, we decided to train the model to predict heads rather than entire spans to optimize the evaluation metric. Having all the useful information (even dependency trees), the model should learn the original rules for selecting the head.
A straightforward approach to get mention heads is to represent mention with its headword on the input. We describe this approach in detail later. However, this method is not ideal because multiple mentions can share the same head, which could result in merging distinct mentions and their clusters. To prevent this, we represent mentions with their full spans, predicting the head of each mention at the top of our model, and then outputting only the headword(s). This way, when building clusters, mentions are represented by their spans, ensuring that clusters of different mentions with the same head are not erroneously merged.
We implemented two versions of the head prediction model, both as separate classification heads on top of our coreference resolution model. The first model predicts the relative position of the headword(s) within a span using the hidden representation of the span from the coreference model. The output probabilities of head positions are obtained using sigmoid activation, allowing the model to predict multiple heads, even though only a single headword is present in the gold data. This serves as an optimization of the evaluation metric: if multiple words are likely headword candidates, it is statistically advantageous to output all of them.
The second model uses a binary classification of each span and head candidate pair, so again, there can be more headwords of a single span predicted.
@@ 4.4 Head Representations
Later, inspired by the word-level coreference resolution model, we decided to try to model mentions only with their syntactic head. As mentioned above, For the CRAC official evaluation metric, reconstructing the original span is unnecessary, as predicting the correct head suffices. Therefore, we explored a simplified word-level coreference resolution approach that bypasses the span prediction step. Given that syntactic information is available in all CorefUD datasets, this is not so unrealistic scenario because in most cases the spans can be reconstructed by simply taking the whole subtree of the predicted node (head). Heads are heuristically selected only in cases where a mention does not form a single subtree. The proportion of such mentions is shown in Table 3(column non-tree). Using a word-level model reduces the mention space from quadratic to linear, making the model more efficient and minimizing potential false-positive mentions. Moreover, we believe that for very long mentions, the standard representation (sum of the start token, end token, and attended sum of all tokens) becomes insufficient.
@@ 4.5 Singletons
Some datasets in the CorefUD collection have singletons annotated and others do not. Specifically, in CorefUD 1.1, 8 out of 17 datasets have more than 10% singletons and 6 of these have more than 70% singletons which is probably a sign of consistent entity annotation independent of the coreference annotation. The baseline model completely ignores singletons during training444The loss is the sum over all correct antecedents and since singletons have no gold antecedents they do not affect the loss.. As a result for these 6 singleton-including datasets, we discard more than 70% of training data for mention identification task. To leverage this data, we incorporate singleton modeling into our model and propose several approaches to address this.
@@ 4.5.1 Another Dummy Antecedent
In the first proposed approach, we introduce another virtual antecedent (with trained embedding) representing that the mention has no real antecedent, but it is a valid mention, so it is a singleton. For these mentions, we avoid using the binary score because it would unnecessarily model the similarity between all the singletons 555because all the singleton representations would be trained to be similar to the dummy antecedent. Additionally, included an option to use a separate feed-forward network for predicting singleton scores (different from the one for other mentions scores).
@@ 4.5.2 Mention Modeling
In the prediction step, the mention score is evaluated only for potential singletons. If a mention has no real antecedent we look at the mention score. If it is likely to be a mention we make it a singleton, otherwise it is not a mention at all.
@@ 4.6 Overlapping Segments
One limitation of the XLM-R model is the short sequence length. In the employed model, the input document is split into segments which are processed individually with a BERT-like encoder and merged in the antecedent-prediction head. To propagate errors correctly across segments, we need to process all segments in a single gradient update. However, due to GPU memory constraints, we set a maximum number of segments, and if a document exceeds this limit, we split it into multiple individual documents without any mutual coreferences.
We propose using segment overlapping with a cluster merging algorithm to address this issue. The cluster merging algorithm is straightforward: We iterate over all mentions in the new part and if it is present in a cluster in the previous part we simply make a union between its clusters from both parts. We experiment with minimal (just one overlapping segment) and maximal (each continuing example has just one new segment the rest is overlapping) segment overlap. For instance, if a document has 6 segments and the maximum number of segments for a single example is 4, then we split the document into 3 examples overlapping with 3 segments. The example is shown in Figure 5. We further extend this method by only using clusters in the continuing examples where at least one mention is present in new segments because if there is no such mention then all the mentions were seen already in the previous part with a longer left context.
@@ 5.1 Training
We trained all models on NVIDIA A40 GPUs using online learning (batch size 1 document). We limit the maximum sequence length to 6 non-overlapping segments of 512 tokens. During training, if the document is longer than tokens, a random segment offset is sampled to take a random continuous block of 6 segments, and the rest of them are discarded. During prediction, longer documents are split into independent sub-documents (for simplicity, non-overlapping again). For the head-only model, we double the maximum number of segments since the model is much more memory-effective therefore we can afford it. We train a model for each dataset for approximately 80k updates in our monolingual experiments. For joined-pre-trained models, we use 80k steps for model pre-training on all the datasets and approximately 30k for fine-tuning on each dataset. Each training took from 8 to 14 hours.
@@ 5.2 Baseline and Extensions
We begin by training a monolingual model for each language. For these experiments, we use XLM-Roberta-large and specific monolingual models for each language. Monolingual models for individual languages are listed in Table 5. Then we train a joined cross-lingual model for all the datasets which we fine-tune for each dataset afterwards.
To evaluate the effect of each extension, we first run all its variants described in 4 on the basic XLM-R-large model and measure the performance for each language (Sections 6.1.1 and 6.1.2). Then we compare each extension in its best configuration to the baseline joined model in Section 6.1.3 (since joined models achieve the best results from our three baseline models). In the last step, we evaluate the effects of Overlapping segments only on the best model configuration for each language (because it is orthogonal to all other extensions).
@@ 5.3 Zero-shot Cross-lingual Experiments
We also conduct zero-shot cross-lingual experiments in two variants; dataset zero shot, and language zero shot. In dataset zero shot we train the model on all the datasets except one and then we evaluate the model on the excluded dataset. However, this approach is not a true cross-lingual zero-shot test, as multiple datasets exist for some languages, meaning the model was still exposed to the language of the evaluation data during training.
To address this, we performed language zero-shot experiments, where we removed all datasets for a particular language from the training data, and then evaluated the model on all datasets in that language. Additionally, we conducted a final experiment focusing on the ParCor corpora, which consists of parallel English and German datasets. This allowed us to perform a true zero-shot evaluation on both ParCor corpora.
@@ 6 Results & Discussion
Table 5 presents the results of baseline models on development data parts. The Monoling column shows the result of the monolingual model specific to each language. The XLM-R column presents results of XLM Roberta large trained separately for each dataset. The Joined column corresponds to the joined model described in Section 4.1. Several key observations can be drawn from these results. Most notably, monolingual models tend to outperform the cross-lingual XLM-Roberta model, especially for the Lithuanian and Czech PCEDT datasets, where the monolingual models—even though they are smaller—exhibit significantly better performance. The reason is probably in the difference of these two datasets from the other ones. We can see that for these two datasets, joined training does not help as much as for other datasets. For the rest of the datasets, joined training surpasses both other models (monolingual and multilingual trained from scratch).
@@ 6.1.1 Span2Head
Table 6 shows the performance of different variants of span2head model. None column shows the results of the baseline XLM-R-large model without any extensions. The Multi column reflects the results of the multi-class relative position classification (the first method described in Section 4.3). The binary column represents the second method (binary classification of span-headword candidate pair). We added Heads column with heads-only mention representation to show if it is better to use the span2head model or heads only model (since they are mutually exclusive). From the results, we can make several conclusions:
@@
Binary model achieves better results, than the multi-class model.
@@
For all the datasets where the span2head model helps, the heads-only model is much better so it does not make sense to use the span2head model in the next experiments.
Another interesting point is that for a monolingual model trained from scratch, the heads-only model improves the results on 13 out of 17 datasets and the results are worse only for a single dataset (English-ParCor).
Dataset/S2H model | None | Multi | Binary | Heads |

### v3

@@ 3.1 Evaluation Metrics
For evaluation, we employ the official metric for CRAC 2023 Multilingual Coreference Resolution shared task implemented via the CoreUD scorer222https://github.com/ufal/corefud-scorer. The metric is a modification of the CoNLL score for coreference resolution, averaging three F1 metrics: CEAF-e, , and MUC. Each of these three metrics has a different universe. MUC operates on the level of coreference links, works with mentions and CEAF-e is entity-based metric. All these three metrics are described later in this section. The main modification of the CorefUD scorer lies in using the head matching. In head matching, the gold and system mentions are considered identical if and only if they have the same syntactic head. Another important aspect is that the primary metric ignores singletons.
3.1.1 MUC Score
@@
The MUC score [16] is a link-based metric. It measures precision and recall of the identification of coreference links in a coreference chain. Coreference chains are formed by linking each mention in a cluster to its antecedent, creating a sequence of references. For instance, precision is computed by summing all correctly predicted links divided by all the links in the system response. In recall, the numerator is the number of all the gold links. Then, we compute the standard F1 measure as the final metric. The number of correctly identified links is technically computed from a partition function , which determines how many entities cover all the mentions of gold entity . Recall is then computed as:
3.1.2
[17] was proposed to address problems of MUC score. One of them is that MUC score as a link-based metric does not address singletons at all. operates on mention level. It computes precision and recall for every mention and then averages these values. the recall and precision of a mention is defined as follows:
3.1.3 CEAF
The problem with is that each entity is addressed many times (with each mention). CEAF [18] tries to solve this problem. It aligns gold and system entities one-to-one, maximizing the similarity of aligned entities. This leads to the maximum bipartite matching problem. Precision and recall are then defined as follows:
where is sum of similarities of all the matching entities in one-to-one mapping found by Kuhn-Munkres Algorithm, and is self-similarity of an entity.
[18] present two entity-pair similarity measures that can be used in CEAF: one is the absolute number of common mentions between two entities, and the other is a “local” mention F-measure between two entities. The two measures lead to the mention-based CEAF-m and entity-based CEAF-e, respectively. [19]
@@ 4 Baseline Model & Extensions
We adopt an end-to-end model from [3], and we use it as the baseline model (without higher-order inference). We use UDAPI333https://github.com/udapi/udapi-python to load the data in CorefUD format. We use XLM-Roberta-large as an encoder.
Building on this baseline, we propose several extensions to the standard end-to-end model presented earlier in this paper. Our objective is to create a universal model suitable for all datasets in the CorefUD collection. We also reevaluate some extensions proposed by [13] for comparison (namely syntax representation and Span2head model).
@@ 4.1 Cross-lingual Training
| mBERT | 180M | 15M |
| XLM-R | 550M | 20M |
Table 4 shows the total number of parameters and the number of parameters trained from scratch. We can see that XLM-R large has approximately 20 million parameters trained from scratch444The number is the average of different configurations since different extensions add different numbers of parameters.. Training so many parameters from scratch can lead to overfitting, especially for smaller datasets. To mitigate this, we first train a joined model on the concatenation of all the datasets and then we finetune this model on each dataset separately.
@@ 4.2 Syntactic Information
We believe that integrating syntactic information can greatly improve the performance. Dependency information is crucial for identifying mention heads, as some datasets in the CorefUD collection label mentions as nodes within the dependency tree rather than spans in the linear text sequence.
We use the syntactic information model from [13]. In this model, we modify a model to encode syntactic information in the following way: For each token, we add its path to ROOT to its embedding representation. To encode a path in a dependency tree, we use a concatenation of BERT embeddings of all the nodes on the path together with the learned embedding of corresponding dependency relations. To have a fixed-length word representation, we set a maximum tree-path length. By adding tree representation, we increase the total size of a token representation in the model by . In our experiments, we use the maximum tree-path length of 5.
@@ 4.3 Span2head Model
Since the official evaluation uses a head-match metric, we trained the model to predict the heads of mentions rather than full spans to better align with the evaluation criterion. With access to all the necessary information, including dependency trees, the model should be able to infer the original rules for selecting heads.
Our goal is to represent mention with a headword only. However, multiple mentions can share the same head. To address this issue, Span2head model represents mentions by their full spans and predicts the head of each mention at the top of the CR model, outputting only the headword(s). This ensures that, during cluster formation, mentions are represented by their spans, preventing incorrect merging of clusters for distinct mentions that share a headword.
Authors of [13] proposed two versions of the head prediction model, both implemented as separate classification layers on top of our coreference resolution model. The first multi-class model predicts the relative position of the headword(s) within a span, using the span’s hidden representation from the coreference model. Head position probabilities are calculated with sigmoid activation, allowing the model to predict multiple headwords, even though only one is present in the gold data. This serves to optimize the evaluation metric by outputting all likely headword candidates. The schema of the model is shown in Figure 3
The second binary model employs binary classification for each span and head candidate pair, which also enables the prediction of multiple headwords for a single span. The schema of the model is shown in Figure 4.
@@ 4.4 Head Representations
Inspired by the word-level coreference resolution model, we decided to try to model mentions only with their syntactic head. As mentioned earlier, this model merges some mentions into one in cases where multiple mentions share the same head. However, according to [4], reduction to headword is highly beneficial. As also mentioned above, for the CRAC official evaluation metric, reconstructing the original span is unnecessary, as predicting the correct head suffices. Therefore, we explored a simplified word-level coreference resolution approach that bypasses the span prediction step. Given that syntactic information is available in all CorefUD datasets, this is not an unrealistic scenario because, in most cases, the spans can be reconstructed by simply taking the whole subtree of the predicted node (head). Heads are heuristically selected only in cases where a mention does not form a single subtree. The proportion of such mentions is shown in Table 3(column non-tree). Using a word-level model reduces the mention space from quadratic to linear, making the model more efficient and minimizing potential false-positive mentions. Moreover, we believe that for very long mentions, the standard representation (sum of the start token, end token, and attended sum of all tokens) becomes insufficient.
@@ 4.5 Singletons
Some datasets in the CorefUD collection have singletons annotated, and others do not. Specifically, in CorefUD 1.1, 8 out of 17 datasets have more than 10% singletons, and 6 of these have more than 70% singletons, which is probably a sign of consistent entity annotation independent of the coreference annotation. The baseline model completely ignores singletons during training555The loss is the sum over all correct antecedents, and since singletons have no gold antecedents, they do not affect the loss.. As a result, for these 6 singleton-including datasets, we discard more than 70% of training data for the mention identification task. To leverage this data, we incorporate singleton modeling into our model and propose several approaches to address this.
@@ 4.5.1 Another Dummy Antecedent
In the first proposed approach, we introduce another virtual antecedent (with trained embedding) representing that the mention has no real antecedent, but it is a valid mention, so it is a singleton. For these mentions, we avoid using the binary score because it would unnecessarily model the similarity between all the singletons 666because all the singleton representations would be trained to be similar to the dummy antecedent. Additionally, we include an option to use a separate feed-forward network for computing singleton scores (different from the one for other mentions scores but trained simultaneously on the top of the same encoder).
@@ 4.5.2 Mention Modeling
In the prediction step, the mention score is evaluated only for potential singletons. If a mention has no real antecedent, we look at the mention score. If it is likely to be a mention, we make it a singleton; otherwise, it is not a mention at all.
@@ 4.6 Overlapping Segments
One limitation of the XLM-R model is the short sequence length. In the employed model, the input document is split into segments, which are processed individually with a BERT-like encoder and merged in the antecedent-prediction head. To propagate errors correctly across segments, we need to process all segments in a single gradient update. However, due to GPU memory constraints, we set a maximum number of segments, and if a document exceeds this limit, we split it into multiple individual documents without any mutual coreferences.
We propose using segment overlapping with a cluster merging algorithm to address this issue. The cluster merging algorithm is straightforward: We iterate over all mentions in the new part, and if it is present in a cluster in the previous part, we simply make a union between its clusters from both parts. We experiment with minimal (just one overlapping segment) and maximal (each continuing example has just one new segment the rest is overlapping) segment overlap. For instance, if a document has 6 segments and the maximum number of segments for a single example is 4, then we split the document into 3 examples overlapping with 3 segments. The example is shown in Figure 5. We further extend this method by only using clusters in the continuing examples where at least one mention is present in new segments. If there is no such mention then all the mentions were seen already in the previous part with a longer left context.
@@ 5.1 Training
We trained our models using NVIDIA A40 GPUs. the batch size was equal to one document. We limit the maximum sequence length to 6 non-overlapping segments of 512 tokens. During training, if a document exceeded this length, a random offset was used to select a continuous block of 6 segments, with the remaining segments discarded. During prediction, longer documents are split into independent sub-documents. In the basic model, there is no overlap between sub-documents. In the case of the head-only model, we increased the number of segments to eight since this model is more memory-efficient, allowing for a larger sequence length. Each model was trained on its respective dataset for around 80,000 updates in our monolingual experiments. For the models pre-trained on multiple datasets, we used 80,000 steps for joint pre-training across all datasets, followed by roughly 30,000 steps for fine-tuning on each individual dataset.
@@ 5.2 Baseline and Extensions
We begin by training a monolingual model for each language. For these experiments, we use XLM-Roberta-large and specific monolingual models for each language. Monolingual models for individual languages are listed in Table 5. Then, we train a joined cross-lingual model for all the datasets, which we fine-tune for each dataset afterward.
To evaluate the effect of each extension, we first run all its variants described in 4 on the basic XLM-R-large model and measure the performance for each language (Sections 6.1.2 and 6.1.3). Then, we compare each extension in its best configuration to the baseline joined model in Section 6.1.5 (since joined models achieve the best results from our three baseline models). In the last step, we evaluate the effects of Overlapping segments only on the best model configuration for each language (because it is orthogonal to all other extensions).
@@ 5.3 Zero-shot Cross-lingual Experiments
We also conduct zero-shot cross-lingual experiments in two variants: dataset zero-shot and language zero-shot. In dataset zero shot, we train the model on all the datasets except one, and then we evaluate the model on the excluded dataset. However, this approach is not a true cross-lingual zero-shot test, as multiple datasets exist for some languages, meaning the model was still exposed to the language of the evaluation data during training.
To address this, we performed language zero-shot experiments, where we removed all datasets for a particular language from the training data. Then, we evaluated the model on all datasets in that language. Additionally, we conducted a final experiment focusing on the ParCor corpora, which consists of parallel English and German datasets. This allowed us to perform a true zero-shot evaluation on both ParCor corpora.
@@ 6 Results & Discussion
Table 5 presents the results of baseline models on development datasets parts. Several key observations can be drawn from these results. Most notably, monolingual models tend to outperform the cross-lingual XLM-Roberta model, especially for the Lithuanian and Czech PCEDT datasets, where the monolingual models—even though they are smaller—exhibit significantly better performance. The reason is probably in the difference of these two datasets from the other ones. We can see that for these two datasets, joined training helps only slightly compared to other datasets. For the rest of the datasets, joined training surpasses both other models (monolingual and multilingual trained from scratch).
@@ 6.1 Effects of Extensions
6.1.1 Joined Training
We further investigate the effect of joined training on the Czech PDT dataset by analyzing concrete documents. We found that the joined model is much better in identifying mentions. For example, if we look at the following part of a document with the gold annotations:
- GOLD
Visegrádskáe2 trojka zřejmě nebude mít volný obchode3. Maďarskoe6 a Polskoe7 se rozhodly ustoupit od chystané zóny volného obchodu mezi zeměmi visegrádskée2 trojky
- GOLD in English
The Visegrade2 Three will probably not have free tradee3. Hungarye6 and Polande7 have decided to withdraw from the planned free trade zone between Visegrade2 Three
- Baseline
Visegrádskáe2 trojka zřejmě nebude mít volný obchod. Maďarskoe3 a Polsko se rozhodly ustoupit od chystané zóny volného obchodu mezi zeměmi Visegrádskée2 trojky.
- Joined
"Visegrádskáe2 trojka zřejmě nebude mít volný obchode3. Maďarskoe6 a Polskoe7 se rozhodly ustoupit od chystané zóny volného obchodu mezi zeměmi visegrádskée2 trojky",
where only the very long entity e4 is missing. (This is caused by maximum mention length, as discussed later).
We have also noticed that the most significant difference between the baseline and joined model is in the case of foreign news, where the same topic can appear in multiple datasets (as this news from the European parliament)
@@ 6.1.2 Span2Head
Table 6 shows the performance of different variants of Span2head model. We added Heads column with heads-only mention representation to show if it is better to use the Span2head model or the heads-only model (since they are mutually exclusive). From the results, we can make several conclusions:
@@
The Binary model achieves better results than the multi-class model.
@@
For all the datasets where the Span2head model helps, the heads-only model is much better, so it does not make sense to use the Span2head model in the next experiments.
Another interesting point is that for a monolingual model trained from scratch, the heads-only model improves the results on 13 out of 17 datasets, and the results are worse only for a single dataset (English-ParCor), which is very small, so the difference might be caused by random chance.
| Dataset/S2H model | None | Multi-class | Binary | Heads |